### v2

@@ 1 Introduction
The rise of foundation models in recent years has led to tremendous developments in understanding natural languages Paa2023 . Although they were originally developed to process and generate written text, these models have transcended their initial purpose due to their generalizable nature and wide applicability. Foundation models have shown great potential in the field of bioinformatics Zhang2023 , since their capacity to be trained on vast amounts of unlabeled data and their adaptability enable them to achieve state-of-the-art performance in a variety of tasks.
Early applications of foundation models in bioinformatics can be seen in analyzing protein sequences intro1 ; intro2 , which were then trained on diverse applications like calculation of protein structure, prediction of mutation effects and the understanding of phylogenetic structure Lupo2022ProteinLM ; Fang2022AMF ; Nijkamp2022ProGen2ET . These models have since evolved beyond proteins into DNA and RNA analysis, and have demonstrated the ability to surpass previous benchmarks in identifying regulatory elements, predicting chromatin profiles, analyzing evolution from genomic sequence data and predicting the impacts of mutations in DNA jiDNABERTPretrainedBidirectional2021 ; DallaTorre2023TheNT ; Nguyen2023HyenaDNALG ; Zvyagin2022GenSLMsGL ; Yamada2021PredictionOR . The ability to visualize and interpret the internal model structure Vig2020BERTologyMB and to derive key insights of the underlying biological processes zhangTransformerGeneExpression2022 demonstrate the unique advantages offered by foundation models in the field of bioinformatics.
@@ 1.1 Limitations of previous work
Prior work on Transformer-based models for DNA sequence analysis exists in two forms: (i) Encoder-only models jiDNABERTPretrainedBidirectional2021 ; fishmanGENALMFamilyOpenSource2023 ; zhangTransformerGeneExpression2022 ; DallaTorre2023TheNT that focus on classification and regression-based downstream tasks and (ii) Decoder-only models Nguyen2023HyenaDNALG ; zhangDNAGPTGeneralizedPretrained2023 that are capable of classification, regression as well as generative tasks that involve design and synthesis.
A combination of encoder and decoder blocks enables the model to perform sequence-to-sequence transformations. One of the fundamental processes undergone by DNA is its transcription into an RNA sequence and subsequent translation into protein sequences, the building blocks of all living organisms. Understanding sequence-to-sequence processes like these is crucial to advancing our knowledge of genetics, and developing an encoder-decoder model is an important step in this direction. Although decoder-only models are capable of sequence-to-sequence transformations, they have no independent means of creating representations of the input sequence, and both input and target tokens are processed in an equivalent fashion. Previous work has shown that a multitask finetuned encoder-decoder Large Language Model (LLM) outperforms decoder-only models on zero-shot generalization, despite being an order of magnitude smaller sanh2022multitask . Since a decoder-only architecture will have a unidirectional framework that attends to the source and target sequence simultaneously, as the length of the target sequence grows, the extent to which the model attends to the source will decrease leading to reduced performance in downstream tasks fu2023decoderonly . Our work demonstrates how the cross-attention layers in the decoder leverage the information in the embeddings generated by the encoder, leading to improved performance in training tasks.
@@
Biological sequences like DNA are encoded using a vocabulary of four symbols (A, C, T, G) representing nucleic acids. These sequences are converted into a Transformer-compatible format by a tokenizer, which generates a list of tokens for any given input. Since these models were initially developed for applications in natural languages, the most prevalent forms of tokenization are sentence-piece or word-piece, where the language vocabulary is built using natural ideas like words or syllables. In the absence of typical indicators of linguistic order in DNA, like spaces and punctuation, these tokenization schemes use statistical techniques to determine the ‘words’ that make up the vocabulary of the input sequences. A few examples of previously used tokenizers are: k-mer jiDNABERTPretrainedBidirectional2021 , SentencePiece DallaTorre2023TheNT , and byte-pair encoding (BPE) fishmanGENALMFamilyOpenSource2023 tokenization. While such techniques identify optimal encoding methods by constructing tokens having multiple base pairs, they are vulnerable to any type of noise present in the sequence. A single variation in a base pair will result in the fragment being mapped to a completely different word in the vocabulary, resulting in an outsized impact from a small perturbation Dotan2023EffectOT . We use a simplified tokenization scheme where each character corresponds to a single token, resulting in a longer average tokenized length, but more resiliance to the variations mentioned above.
@@ 1.2 Our contributions
We evaluate the performance of the ENBED foundation model on sequence-level classification tasks and compare it’s accuracy against contemporary foundation models. We show that ENBED outperforms the state-of-the-art in 22 of the 25 benchmarks devised by the authors of the Nucleotide Transformer DallaTorre2023TheNT and Genomic Benchmarks Greov2022GenomicBA datasets. These benchmarks consist of tasks like identifying enhancers, promotors, splice sites and histone marks in multi-species data comprising of genomic sequences from human, mouse, yeast, fruit fly and worm DNA.
@@
Long-read sequencing using Nanopores is used to study telomeres, which are protective caps found at chromosomal ends and have long repetitive elements. It has been found that telomeres in many organisms are frequently miscalled Tan2022 , referring to errors in the process that translates electrical signals into the alphabet of DNA. We illustrate how ENBED can focus on fragments that look incorrect or out of place, demonstrating the model’s ability of distinguishing between noisy and accurate data. In a synthetic dataset constructed using noise distributions found in real-world raw sequence data, we demonstrate that our model can identify sequences containing noise with an accuracy of 97.1%, leveraging the information internalized by bring pretrained on the telomere-to-telomere reference sequences.
@@
Mapping the complete human genome was a significant milestone in modern biology, and it has produced a new set of challenges in identifying the functions and interactions of different parts of the genome. We fine-tune our model to solve a version of this problem by identifying the biological functions of genomic sequences among the nine most common functional classes using a fine-tuned model, achieving an $F_{1}$ score of 74.1.
@@ 2.1 Encoder-Decoder Model Architecture
Figure 1: Model Architecture. The model is constructed using encoder and decoder blocks with a ratio of 2:1. Both types of blocks consist of attention and feed-forward layers, with the decoder blocks additionally incorporating the embeddings in encoder-decoder attention layers. (Sec 2.1)
ENBED is built using an encoder-decoder architecture (Fig. 1) consisting of encoder and decoder blocks, each comprised of two subcomponents: an attention layer and a feed-forward neural network. The attention layers process a sequence by replacing each element with a weighted average of the rest of the sequence, after which they are normalized and passed through the feed-forward neural network. Dropout is applied to the feed-forward network, the attention weights, and the input and output of the entire stack. The implementation is written using JAX jax2018github and the Flax-former library flax2020github .
We formulate a model with 1.2B trainable parameters, with the configuration specified in Table 6. The model is encoder-heavy since idiosyncratic relationships among tokens are better encoded by devoting a larger share of parameters to these blocks. We found that adjusting the encoder-to-decoder ratio to 2:1 improved performance, with a  1% increase in Masked Language Modeling (MLM) accuracy for all model sizes over the 3:1 ratio chosen by the authors of ByT5 xueByT5TokenfreeFuture2022 , a similar architecture built to process token-free text-to-text transformations. We also find that reducing the masked span length, which is the average number of tokens masked during pre-training, from 40 down to 20 helps in faster convergence owing to the significantly smaller vocabulary of DNA.
@@ 2.3 Attention
Attention can be understood as a soft-lookup of a query Q in a dictionary of stored keys K and values V. Attention scores are generated by calculating the similarity between Q and K, each having a dimension $d$, with scaled dot-product attention $\left(\text{Softmax}\left(\textbf{Q}\textbf{K}^{T}/\sqrt{d}\right)V\right)$ being the most common implementation. Increasing the sequence length $L$ can be a challenge, since this type of attention has a complexity of $O(L^{2})$.
@@ 2.5 Application Domains
The Genomic Benchmarks (GB) dataset consists of sequences from four organisms: Human, mouse (Mus musculus), roundworm (Caenorhabditis elegans) and fruit fly (Drosophila melanogaster). The dataset comprises of: (i) Human enhancers from Cohn et. al. Cohn2018EnhancerIU and Ensembl Martin2022Ensembl2 , (ii) Open Chromatin Region classifications from the Ensembl build, (iii) Computationally generated data for coding and non-coding sequences (iv) Multi-class data composed of three regulatory elements (promotors, enhancers and Open Chromatin Regions), (v) Non-TATA promotor sequences imported from Umarov et. al. umarov2017recognition .
@@
The Nucleotide Transformer (NT) benchmarks consist of five data sources: (i) Epigenetic marks in the yeast genome, which use experimentally obtained nucleosome occupancy values processed into positive and negative observations and to provide the following histone marks datasets: {H3, H4, H3K9ac, H3K14ac, H4ac, H3K4me1, H3K4me2, H3K4me3, H3K36me3, and H3K79me3}, (ii) A dataset geng2022deep consisting of a mix of strong, weak and non-enhancers. (iii) Promotor sequences $300$ base pairs in length around transcription start sites, divided on the basis of TATA and non-TATA box promotors. (iv) Splice site datasets composed of donor, acceptor and non-splice site sequences from phylogenetically diverse organisms.
@@
We generate a synthetic dataset with segments of $512$ nucleotides selected at random from TeloBase Lycka2023 , a comprehensive database of information about telomere motif diversity. We introduce noise based on real-world raw DNA sequencing data to generate negative samples. The process introduces insertion, deletion and base-call mismatch errors with a uniform distribution of 2% each, resulting in a simulated read accuracy of 94%. The model is fine-tuned on a sequence classification task with this labeled dataset. This process can be likened to out-of-distribution detection Fort2021ExploringTL , since the negative samples would represent data that does not belong to the distribution of the training dataset.
@@
We can formulate the process of annotating genes as a sequence-to-sequence task, with the input being a DNA sequence fragment and the output containing information regarding the name, type and function of the fragment. For evaluating our model, we train it to output the biological function annotation of a given genomic input sequence up to 512 base pairs in length. We choose the following annotation types for our experiment: Coding Sequences, IncRNA, snoRNA, miscRNA, miRNA, snRNA, TEC, Processed and Unprocessed Pseudogenes. These annotations are obtained from the Ensembl dataset Martin2022Ensembl2 .
@@
Figure 2: Phylogenetic Tree. Circular cladogram visualization of the Influzenza H1 gene sequences. Nodes are represented by yellow dots.
Human influenza A viruses are named based on the geographic location where the virus was isolated, the date of the isolate, and the identity of the two major surface proteins, hemagglutinin (HA) and neuraminidase (NA). We choose the HA1 sequences to create the Influenza virus mutation dataset, selecting the segments with most highly variable regions for training and validation. We obtain our source data from Berman2020MutaGANAS and subset the HA1 nucleotide sequence of the H3N2 Influenza virus between 300 to 799 bp (100-266 amino-acids) to capture the Antigenic site A and B. The selected region is a part of the globular domain that occurs in a jelly-roll fold of eight-stranded anti-parallel beta-sheets, containing the most commonly mutating amino-acid residues around the receptor binding site. The HA1 head also accumulates N-linked glycosylation sites over time, which are thought to mask antigenic sites from immune recognition. The glycosylation of the HA1 globular domain modulates receptor binding, stimulates host antibody responses, and shields key antigenic sites to facilitate immune evasion of the virus. By focusing on the HA1 subdomain, we aimed to evaluate the sequence-to-sequence model on a functionally important region of influenza HA that experiences significant antigenic drift and glycosylation changes. We visualize the H1 gene sequences using a circular cladogram (Fig. 2).
@@ 3 Results
Upon convergence, the pre-training process yields a foundation model ready to be applied to downstream tasks. The initial layers in the pre-trained model are frozen since they contain generalizable information that helps the model build versatile internal representations of the data. We visualize these internal representations by extracting the encoder output layer and plotting attention maps in Fig. 3.
@@
Figure 3: Interpreting Attention Layers. We visualize the 12 attention heads of the pre-trained ENBED (Base) foundation model. These maps are generated using the outputs from the final encoder block. The use of multiple attention heads grants the model the ability to simultaneously use a diverse range of patterns to analyze input sequences. We observe that some heads are dedicated to analyzing close neighbors (3, 9, 10) while others display a more dilated version of this phenomenon (1, 2, 5, 11). Additionally, there are heads which attempt to exclude local information and focus on a more global view of the input sequence (4, 6, 8, 12).
3.1 ENBED outperforms state-of-the-art on genomic benchmark datasets

### v3

@@ 1 Introduction
The rise of foundation models in recent years has led to tremendous developments in understanding natural languages Paaß and Giesselbach (2023). Although they were originally developed to process and generate written text, these models have transcended their initial purpose due to their generalizable nature and wide applicability. Foundation models have shown great potential in the field of bioinformatics Zhang et al. (2023b), since their capacity to be trained on vast amounts of unlabeled data and their adaptability enable them to achieve state-of-the-art performance in a variety of tasks.
Early applications of foundation models in bioinformatics can be seen in analyzing protein sequences Elnaggar et al. (2022); Rives et al. (2021), which were then trained on diverse applications like calculation of protein structure, prediction of mutation effects and the understanding of phylogenetic structure Lupo et al. (2022); Fang et al. (2022); Nijkamp et al. (2022). These models have since evolved beyond proteins into DNA and RNA analysis, and have demonstrated the ability to surpass previous benchmarks in identifying regulatory elements, predicting chromatin profiles, analyzing evolution from genomic sequence data and predicting the impacts of mutations in DNA Ji et al. (2021); Dalla-Torre et al. (2023); Nguyen et al. (2023); Zvyagin et al. (2022); Yamada and Hamada (2021). The ability to visualize and interpret the internal model structure Vig et al. (2020) and to derive key insights of the underlying biological processes Zhang et al. (2022) demonstrate the unique advantages offered by foundation models in the field of bioinformatics.
@@ 1.1 Limitations of previous work
Prior work on Transformer-based models for DNA sequence analysis exists in two forms: (i) Encoder-only models Ji et al. (2021); Fishman et al. (2023); Zhang et al. (2022); Dalla-Torre et al. (2023) that focus on classification and regression-based downstream tasks and (ii) Decoder-only models Nguyen et al. (2023); Zhang et al. (2023a) that are capable of classification, regression as well as generative tasks that involve design and synthesis.
A combination of encoder and decoder blocks enables the model to perform sequence-to-sequence transformations. One of the fundamental processes undergone by DNA is its transcription into an RNA sequence and subsequent translation into protein sequences, the building blocks of all living organisms. Understanding sequence-to-sequence processes like these is crucial to advancing our knowledge of genetics, and developing an encoder-decoder model is an important step in this direction. Although decoder-only models are capable of sequence-to-sequence transformations, they have no independent means of creating representations of the input sequence, and both input and target tokens are processed in an equivalent fashion. Previous work has shown that a multitask finetuned encoder-decoder Large Language Model (LLM) outperforms decoder-only models on zero-shot generalization Sanh et. al. (2022) as well as targeted tasks like machine translation Raffel et al. (2020); Fu et al. (2023). Since a decoder-only architecture will have a unidirectional framework that attends to the source and target sequence simultaneously, as the length of the target sequence grows, the extent to which the model attends to the source will decrease leading to reduced performance in downstream tasks Fu et al. (2023). Our work demonstrates how the cross-attention layers in the decoder leverage the information in the embeddings generated by the encoder, leading to improved performance in training tasks.
@@
Biological sequences like DNA are encoded using a vocabulary of four symbols (A, C, T, G) representing nucleic acids. These sequences are converted into a Transformer-compatible format by a tokenizer, which generates a list of tokens for any given input. Since these models were initially developed for applications in natural languages, the most prevalent forms of tokenization are sentence-piece or word-piece, where the language vocabulary is built using natural ideas like words or syllables. In the absence of typical indicators of linguistic order in DNA, like spaces and punctuation, these tokenization schemes use statistical techniques to determine the ‘words’ that make up the vocabulary of the input sequences. A few examples of previously used tokenizers are: k-mer Ji et al. (2021), SentencePiece Dalla-Torre et al. (2023), and byte-pair encoding (BPE) Fishman et al. (2023) tokenization. While such techniques identify optimal encoding methods by constructing tokens having multiple base pairs, they are vulnerable to any type of noise present in the sequence. A single variation in a base pair will result in the fragment being mapped to a completely different word in the vocabulary, resulting in an outsized impact from a small perturbation Dotan et al. (2023). We use a simplified tokenization scheme where each character corresponds to a single token, resulting in a longer average tokenized length, but more resiliance to the variations mentioned above.
@@ 1.2 Our contributions
We evaluate the performance of the ENBED foundation model on sequence-level classification tasks and compare it’s accuracy against contemporary foundation models. We show that ENBED outperforms the state-of-the-art in 21 of the 25 benchmarks devised by the authors of the Nucleotide Transformer Dalla-Torre et al. (2023) and Genomic Benchmarks Grevsova et al. (2022) datasets. These benchmarks consist of tasks like identifying enhancers, promotors, splice sites and histone marks in multi-species data comprising of genomic sequences from human, mouse, yeast, fruit fly and worm DNA.
@@
Long-read sequencing using Nanopores is used to study telomeres, which are protective caps found at chromosomal ends and have long repetitive elements. It has been found that telomeres in many organisms are frequently miscalled Tan et al. (2022), referring to errors in the process that translates electrical signals into the alphabet of DNA. We illustrate how ENBED can focus on fragments that look incorrect or out of place, demonstrating the model’s ability of distinguishing between noisy and accurate data. In a synthetic dataset constructed using noise distributions found in real-world raw sequence data, we demonstrate that our model can identify sequences containing noise with an accuracy of 97.6%, leveraging the information internalized by bring pretrained on the telomere-to-telomere reference sequences.
@@
Mapping the complete human genome was a significant milestone in modern biology, and it has produced a new set of challenges in identifying the functions and interactions of different parts of the genome. We fine-tune our model to solve a version of this problem by identifying the biological functions of genomic sequences among the most common functional classes using a fine-tuned model, achieving an $F_{1}$ score of 74.1.
@@ 2.1 Encoder-Decoder Model Architecture
Figure 1: Model Architecture. The model is constructed using encoder and decoder blocks with a ratio of 2:1. Both types of blocks consist of attention and feed-forward layers, with the decoder blocks additionally incorporating the embeddings in encoder-decoder attention layers.
ENBED is built using an encoder-decoder architecture (Fig. 1) consisting of encoder and decoder blocks, each comprised of two subcomponents: an attention layer and a feed-forward neural network. The attention layers process a sequence by replacing each element with a weighted sum of linear transformations of the input embeddings, after which they are normalized and passed through the feed-forward neural network. Dropout is applied to the feed-forward network, the attention weights, and the input and output of the entire stack. The implementation is written using JAX Bradbury et al. (2018) and the Flax-former library Heek et al. (2023).
We formulate a model with 1.2B trainable parameters, with the configuration specified in Table 6. The model is encoder-heavy since idiosyncratic relationships among tokens are better encoded by devoting a larger share of parameters to these blocks. We found that adjusting the encoder-to-decoder ratio to 2:1 improved performance, with a  1% increase in Masked Language Modeling (MLM) accuracy for all model sizes over the 3:1 ratio chosen by the authors of ByT5 Xue et al. (2022), a similar architecture built to process token-free text-to-text transformations. We also find that reducing the masked span length, which is the average number of tokens masked during pre-training, from 40 down to 20 helps in faster convergence owing to the significantly smaller vocabulary of DNA.
@@ 2.3 Attention
Attention can be understood as a soft-lookup of a query Q in a dictionary of stored keys K and values V. Attention scores are generated by calculating the similarity between Q and K, each having a dimension $d$, with scaled dot-product attention $\left(\text{Softmax}\left(\textbf{Q}\textbf{K}^{T}/\sqrt{d}\right)V\right)$ being the most common implementation. Increasing the sequence length $L$ can be a challenge, since this type of attention has a complexity of $O(L^{2})$. This sets a limit of $L\leq 512$ tokens on our hardware (NVIDIA A100 (40 GB) GPUs).
@@
The aggregated blocks constructed in this procedure resemble previous tokenization schemes like k-mer, used by previous models like DNABERT Ji et al. (2021) and BPE used by GENA-LM Fishman et al. (2023). Our method uses a combination of these aggregated blocks along with higher-granularity local context to achieve a balance between the two, allowing us to process sequences with greater precision.
@@ 2.5 Application Domains
The ENBED foundation model is evaluated across a set of genomic analysis tasks to demonstrate its versatility and the unique advantages of its encoder-decoder architecture. We begin with the Genomic Benchmarks and Nucleotide Transformer Benchmarks, which provide standardized comparisons against existing models for fundamental sequence classification tasks. The noise identification task assesses ENBED’s ability to distinguish genuine sequences from artifacts, leveraging its byte-level precision. Biological function annotation tests the model’s capacity to associate sequence patterns with higher-level functions, crucial for genome interpretation. Finally, the mutation generation task is an end-to-end evaluation of the ENBED, a novel architecture not present in previous genomic language models. This sequence-to-sequence task, focused on predicting viral mutations, showcases ENBED’s potential for modeling complex genomic transformations.
@@
The Genomic Benchmarks (GB) dataset consists of sequences from four organisms: Human, mouse (Mus musculus), roundworm (Caenorhabditis elegans) and fruit fly (Drosophila melanogaster). The dataset comprises of: (i) Human enhancers from Cohn et. al. Cohn et al. (2018) and Ensembl Martin et. al. (2022), (ii) Open Chromatin Region classifications from the Ensembl build, (iii) Computationally generated data for coding and non-coding sequences (iv) Multi-class data composed of three regulatory elements (promotors, enhancers and Open Chromatin Regions), (v) Non-TATA promotor sequences imported from Umarov et. al. Umarov and Solovyev (2017).
@@
The Nucleotide Transformer (NT) benchmarks consist of five data sources: (i) Epigenetic marks in the yeast genome, which use experimentally obtained nucleosome occupancy values processed into positive and negative observations and to provide the following histone marks datasets: {H3, H4, H3K9ac, H3K14ac, H4ac, H3K4me1, H3K4me2, H3K4me3, H3K36me3, and H3K79me3}, (ii) A dataset Geng et al. (2022) consisting of a mix of strong, weak and non-enhancers. (iii) Promotor sequences $300$ base pairs in length around transcription start sites, divided on the basis of TATA and non-TATA box promotors. (iv) Splice site datasets composed of donor, acceptor and non-splice site sequences from phylogenetically diverse organisms.
@@
We generate a synthetic dataset with segments of $512$ nucleotides selected at random from TeloBase Lyčka et al. (2023), a comprehensive database of information about telomere motif diversity. We introduce noise based on real-world raw DNA sequencing data to generate negative samples. Previous work Rabadan et al. (2017) finds that noise in sufficiently deep DNA sequencing data can be approximated by aggregating negative binomial distributions. Using this method, we create a balanced dataset with positive and negative samples. The model is fine-tuned on a sequence classification task with this labeled dataset. This process can be likened to out-of-distribution detection Fort et al. (2021), since the negative samples would represent data that does not belong to the distribution of the training dataset. We describe this procedure in more detail in the Supplementary Material (Section B).
@@
We can formulate the process of annotating genes as a classification task, with the input being a DNA sequence fragment and the output being the class probabilities for the annotation types defined below. For evaluating our model, we train it to output the biological function annotation of a given genomic input sequence up to 512 base pairs in length. We choose the following annotation types for our experiment: Coding Sequences, IncRNA, snoRNA, miscRNA, miRNA, snRNA, TEC, Processed and Unprocessed Pseudogenes. These annotations are obtained from the Ensembl dataset Martin et. al. (2022), and the constructed dataset has an equal number of examples for all classes. We generate 9216 training examples and 1024 validation examples for this task.
@@
Human influenza A viruses are named based on the geographic location where the virus was isolated, the date of the isolate, and the identity of the two major surface proteins, hemagglutinin (HA) and neuraminidase (NA). We choose the HA1 sequences to create the Influenza virus mutation dataset, selecting the segments with most highly variable regions for training and validation. We obtain our source data from Berman et al. (2020) and subset the HA1 nucleotide sequence of the H3N2 Influenza virus between 300 to 799 bp (100-266 amino-acids) to capture the Antigenic site A and B. The selected region is a part of the globular domain that occurs in a jelly-roll fold of eight-stranded anti-parallel beta-sheets, containing the most commonly mutating amino-acid residues around the receptor binding site. The HA1 head also accumulates N-linked glycosylation sites over time, which are thought to mask antigenic sites from immune recognition. The glycosylation of the HA1 globular domain modulates receptor binding, stimulates host antibody responses, and shields key antigenic sites to facilitate immune evasion of the virus. By focusing on the HA1 subdomain, we aimed to evaluate the sequence-to-sequence model on a functionally important region of influenza HA that experiences significant antigenic drift and glycosylation changes. The Supplementary Material contains additional details about the construction of training and validation splits for the dataset.
@@ 3 Results
Upon convergence, the pre-training process yields a foundation model ready to be applied to downstream tasks. The initial layers in the pre-trained model are frozen since they contain generalizable information that helps the model build versatile internal representations of the data. We visualize these internal representations by extracting the encoder output layer and plotting attention maps in Fig. 2. These maps are generated using the outputs from the final encoder block. The use of multiple attention heads grants the model the ability to simultaneously use a diverse range of patterns to analyze input sequences. In Fig 2, we observe that some heads are dedicated to analyzing close neighbors (3, 9, 10) while others display a more dilated version of this phenomenon (1, 2, 5, 11). Additionally, there are heads which attempt to exclude local information and focus on a more global view of the input sequence (4, 6, 8, 12).
Figure 2: Interpreting Attention Layers. We visualize the twelve attention heads of the pre-trained ENBED foundation model.
3.1 ENBED outperforms state-of-the-art models on genomic benchmark datasets